OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation −9 + 8v = 7 + 4v is given below. Choose the step that is missing from the solution. −9 + 8v = 7 + 4v ( ) 4v = 16 v = 4 ----------------------- 8v = −2 + 4v −9 + 12v = 7 8v = 16 + 4v −2 + 8v = 4v

OpenStudy (anonymous):

are these all different questions?

OpenStudy (anonymous):

ok, first you need to combine like terms and simply on the first one, the step would be 8v=7+4v+9.....4v=16

OpenStudy (anonymous):

or 8v=16+4v